Obtain an expression for the entropy of an assembly of photons. Using this expression, show that the equation for the adiabate of a photon gas is given by pV4/3= constant.
1
Expert's answer
2020-06-16T10:24:57-0400

Enthalpy : "H = U+P\times V"


"S = \frac {4}{3} (\frac{U}{T})"


"S = \frac {4}{3}T^3 \alpha V = constraint"


Thus "V\times T^3= C"


since "P = \frac{1}{3}T^4 \alpha V , V\times T^3 = V\times P^ { \frac {3}{4}} or PV^{\frac {4}{3}}=C."